# ATI-CTLO: Adaptive Temporal Interval-based Continuous-Time LiDAR-Only Odometry

### v2

@@ Abstract
The motion distortion in LiDAR scans caused by the robot’s aggressive motion and environmental terrain feature significantly impacts the positioning and mapping performance of 3D LiDAR odometry. Existing distortion correction solutions struggle to balance computational complexity and accuracy. In this letter, we propose an Adaptive Temporal Interval-based Continuous-Time LiDAR-only Odometry, which based on straightforward and efficient linear interpolation. Our method can flexibly adjust the temporal intervals between control nodes according to the dynamics of motion and environmental degeneracy. This adaptability enhances performance across various motion states and improves the algorithm’s robustness in degenerate, particularly feature-sparse, environments. We validated our method’s effectiveness on multiple datasets across different platforms, achieving comparable accuracy to state-of-the-art LiDAR-only odometry methods. Notably, in situations involving aggressive motion and sparse features, our method outperforms existing LiDAR-only methods.
@@ I INTRODUCTION
LiDAR odometry, recognized for its precise measurement capabilities and adaptability to changing light and weather conditions [1], has been extensively utilized in robotics and autonomous driving, showing remarkable progress over the last decade. Nonetheless, LiDAR odometry faces significant practical challenges, particularly the motion distortion in point clouds, which crucially affects localization accuracy and map-building quality. There are several sources of motion distortion, including robot motion during data acquisition, uneven terrain that causes the robot to oscillate, and environmental disturbances such as air currents affecting unmanned aerial vehicles (UAVs). Fig. 1 shows the motion distortion caused by aggressive turn during data acquisition and the mapping result of our method.
@@
Most current LiDAR odometry methods are based on discrete-time methods, which assume LiDAR remains stationary during data acquisition and handle LiDAR distortion as part of preprocessing, aligning all points in a point cloud to a common reference frame based on constant velocity assumption [2, 3]. This method, while straightforward and efficient, is inadequate for more aggressive motions (As Fig. 1b shows,  [3] still remains distortion). In LiDAR-inertial odometry (LIO) systems, IMUs are used to directly measure the poses of LiDAR points to eliminate distortions [4, 5]. However, this requires precise calibration and is highly susceptible to platform vibrations, such as mechanical vibrations of quadruped robots [6], and is limited by the IMU’s measurement range in extreme motion conditions [7].
@@
Figure 1: Motion distortion correction. (a): A single frame of point cloud is distorted due to aggressive turning. (b): Correction distortion under constant velocity assumption (The mapping result of KISS-ICP [3]). (c): Our method, utilizing dynamic temporal interval control, ensures more accurate mapping.
@@
In contrast, continuous-time methods consider the LiDAR to be moving during data acquisition and model the trajectory as a function of time, allowing for pose queries at any given time, offering a new approach to correcting motion distortion. Current models often assume constant velocity [8, 9, 10] or higher-order modeling [11, 12, 13]. However, higher-order models increase computational complexity, while lower-order models suffer from reduced precision. One solution to this problem relies on the assumption of motion continuity, using lower-order modeling but with smaller temporal intervals between the control nodes [8, 14]. Small temporal interval allows the constant velocity motion prior to accommodate aggressive movements, but also introduces another problem. The small temporal intervals in continuous-time methods are often shorter than the duration of a single LiDAR frame, making it essential to segment the point cloud temporally. This segmentation reduces constraints between control nodes, leading to potential algorithm degeneracy or failure in feature-sparse environments.
@@
To address aforementioned problems, we propose a continuous-time LiDAR-only method, the main contributions of our method are as follows:
@@
  • To address the issue of degeneracy due to cloud segmentation, we introduce a degeneracy management method for continuous-time LiDAR odometry. By categorizing level of degeneracy based on the strength of environmental information constraints on the pose, we implement targeted treatments for each level. This approach significantly enhances the robustness of odometry in degenerate environments.
  • The effectiveness of our method has been validated across datasets from multiple platforms. To our knowledge, this is the only LiDAR-only odometry approach that has been experimentally validated on Unmanned Ground Vehicles (UGVs), UAVs, and quadruped robots. The results shows a 10% improvement in positioning accuracy over the state-of-the-art (SOAT) methods. Moreover, in challenging environments, particularly those with sparse features, our algorithm outperforms all existing LiDAR-only odometry methods.
@@ II-A Discrete-time and Continuous-time LiDAR Odometry
3D LiDAR odometry technology primarily estimates pose through point cloud registration. The classic method, Iterative Closest Point (ICP) algorithm [15], involves finding matching point pairs between the target and source point clouds and optimizing for the best transformation to minimize the distance between all pairs. In recent years, several outstanding works have enhanced ICP performance. For instance, KISS-ICP [3] proposed a lightweight ICP method that dynamically adjusts the data association threshold, adapting to multiple scenarios with just one parameter file. Feature-based methods achieve rapid and precise registration by extracting significant points from the point cloud. LOAM [16] is a representative of this approach, inspiring subsequent advancements like LeGO-LOAM [17], which incorporated ground optimization, and F-LOAM [2], which increased efficiency without sacrificing accuracy by adopting a scan-to-map method.
However, these methods rely on discrete-time representation, which inherently leads to LiDAR distortion issues. The common solutions involve using a constant velocity assumption or leveraging IMU pose estimation [4, 5]. Continuous-time trajectory methods offer an alternative approach to removing motion distortions in point clouds, which allow to query poses at any given time. Over the past decade, these techniques have made significant advancements in real-time performance and accuracy. Continuous-time methods can be divided into two main categories: nonparametric and parametric.
The nonparametric methods view the continuous-time trajectory as Gaussian process [18, 19, 20], the GP prior smooths terms between discrete-time nodes. For parametric methods, B-spline is widely utilized to model the trajectory due to the locality and smoothness characteristics[21].Thanks to the efficient derivative computation of cumulative B-splines[22], numerous high-performing B-spline-based continuous-time LiDAR(-IMU) odometry[11, 23] approaches have been developed. These approaches excel in handling point cloud motion distortion and improving the smoothness of the trajectory.
@@ II-B Degeneracy Detection
LiDAR odometry is highly susceptible to degeneracy in environments with sparse or repetitive geometric features. Continuous-time LiDAR odometry, which uses smaller time intervals for more accurate trajectory fitting, exacerbates the risk of degeneracy due to the reduced number of LiDAR points between adjacent control nodes. Hence, degeneracy detecting is essential for improving the robustness of LiDAR odometry.
Zhang et al. [25] determined degeneracy by comparing the smallest eigenvalue $\lambda_{min}$ of the Hessian matrix to a predefined threshold, separating the degenerate directions in the state space and optimizing only in well-constrained directions. Hinduja et al. [26] applied this method to ICP, and using the ratio between the largest and smallest eigenvalues as the eigenvalue threshold. Ren et al. [27] suggested that pose updates in the degenerate direction cause oscillations and calculates the covariance matrix of pose updates during optimization, summing the diagonal elements as the degeneracy factor. Nubert et al. [28] introduced a learning-based degeneracy detection method, directly estimating localizability on raw LiDAR scans.
Another approach directly identifies the degenerate directions through the geometric characteristics of environment. Yang et al. [29] utilized PCA to analyze the normal vectors of all points in the point cloud, also using the smallest eigenvalue $\lambda_{min}$ as a criterion for describing degeneracy. Zhen et al. [30] evaluated the strength of the constraint by measuring the sensitivity of measurements w.r.t the robot pose, but the method requires a prior-build map of the environment. X-ICP [31] employed a similar method to [30], evaluated the contribution of information pairs to the constraints. The method categorized localizability into three levels {localizable, partially localizable, nonlocalizable} and enhanced constraints for partial localizability through resampling the available information.
@@ III-A System Overview
Figure 2: System Overview. Firstly, LiDAR scans segmentation is based on PCA assessment, followed by solving through sliding window approach. Simultaneously, degeneracy management merges cloud segments when degeneracy occurs.
The Fig.2 shows our pipeline. We denote the trajectory as $\mathbf{T}(t)$ in world frame, which is composed of continuous linear segments $[\mathbf{T}_{i},\mathbf{T}_{i+1})$ over interval $[t_{i},t_{i+1})$, the Lie Group $\mathbf{T}_{i}$ represents the pose at timestamp $t_{i}$. For the latest LiDAR input point cloud $\mathbf{L}_{curr}$, PCA technique is first used to evaluate the angle $\boldsymbol{\varphi}$ between the principal directions of $\mathbf{L}_{curr}$ and the previous point cloud $\mathbf{L}_{pre}$. Then the temporal interval $\Delta t$ between control nodes is reduced by a certain factor $\alpha$ once the angle $\boldsymbol{\varphi}$ exceeds the predefined threshold (Section III-B). After setting the temporal interval $\Delta t$, the points in $\mathbf{L}_{curr}$ are divided into cloud segments $\{\mathbf{S}_{i+1},\mathbf{S}_{i+2},\cdots,\mathbf{S}_{i+5}\}$ according to their timestamps. A sliding window approach is utilized to optimize multiple segments $\{\mathbf{S}_{i},\mathbf{S}_{i+1},\mathbf{S}_{i+2}\}$ at once, Once the optimization is finished, the slide window shifts backward to include new linear segments, while the oldest segment is added to the local map (Section III-C). Notably, during the iterative optimization, the degeneracy detection module is applied to determine if the time interval should be increased, which is then achieved by merging linear segments (Section III-D).
@@ III-B PCA Pre-assessment to Reduce Temporal Interval
Figure 3: PCA pre-assessment. A sudden change in the eigenvectors $\boldsymbol{v}_{i}$ indicates a aggressive turn in motion. In the upper image, the nodes with numbers show significant changes in the PCA direction. However, only the yellow nodes (2, 3, 4) are due to motion turns, while the red nodes (1, 5) result from environmental changes (e.g., the robot moving from a room into a hallway, or vice versa). The line chart below records the maximum PCA direction change $\boldsymbol{\varphi}_{max}$ (max_varphi, blue) and the maximum eigenvalue change $\mathcal{V}_{max}$ (max_V, yellow) throughout the entire trajectory. At nodes 1 and 5, there is a significantly change in $\mathcal{V}_{max}$, and we filter out erroneous estimates (red nodes) by setting a threshold(red line).
@@
Inspired by the scan-to-scan principle in [16], we use the matching of the current and previous LiDAR frames to acquire pose transformations. However, differing from the discrete-time method, our scan-to-scan method requires only a rough description of motion changes, not the exact pose transformation between frames, since the kinematic constraints have provided pose priors. Principal Component Analysis (PCA) is used to extract the main direction of the point cloud, with changes in this principal direction indicating the degree of pose change. As shown in Fig. 3, during the robot’s turning process (yellow nodes 2, 3, 4), the principal direction of the point cloud undergoes significant changes (blue). Note that we only concentrate on the changes in the principal direction caused by violent rotations, since constant velocity prior introduces significant errors when such rotations occur.
For the $j$-th point ${}^{L}\mathbf{p}_{j}$ in $N$ points in $\mathbf{L}_{curr}$, we first compute the covariance matrix $\mathbf{C}\in\mathbb{R}^{3\times 3}$ of all points:
$\mathbf{C}$ $=\sum_{j=1}^{N}(^{L}\mathbf{p}_{j}-\overline{\mathbf{p}})^{\top}(^{L}\mathbf{p}_{j}-\overline{\mathbf{p}})$ (1a)
$\overline{\mathbf{p}}$ $=\frac{1}{N}\sum_{j=1}^{N}{{}^{L}\mathbf{p}_{j}}$ (1b)
Then Singular Value Decomposition (SVD) is applied to $\mathbf{C}$ to derive the eigenvalues $\lambda_{0}<\lambda_{1}<\lambda_{2}$ and corresponding eigenvectors $\boldsymbol{v}_{0},\boldsymbol{v}_{1},\boldsymbol{v}_{2}$. The eigenvectors represent the principle directions of LiDAR scans, so the angles $\boldsymbol{\varphi}=[\boldsymbol{\varphi}_{1},\boldsymbol{\varphi}_{2},\boldsymbol{\varphi}_{3}]$ between the principle directions $\boldsymbol{v}^{curr},\boldsymbol{v}^{last}$ of adjacent LiDAR scans indicate the dynamics of motion.
@@
However, the principal directions derived from PCA can abruptly change due to changes in the environment. During our experiments, we observed that these environmental changes also lead to variations in the eigenvalues, which can been seen in Fig. 3. Leveraging this observation, we filter out changes in the PCA principal direction resulting from environmental changes:
@@
Only when the max change $\mathcal{V}_{max}$ in eigenvalues falls within a specified threshold $\mathcal{K}_{val}$ do we consider the principal direction change to be caused by motion. We analyze the final $\boldsymbol{\varphi}_{final}$ by comparing the max element $\boldsymbol{\varphi}_{max}$ with the certain threshold $\mathcal{K}_{vec}$. If $\boldsymbol{\varphi}_{max}$ exceeds $\mathcal{K}_{vec}$ , we reduce the temporal interval by a predetermined factor $\alpha$.
III-C Non-Uniform Linear Interpolation Continuous-time LiDAR-Only Odometry baesed on Sliding Window
The continuous-time LiDAR odometry addresses the following problem: given the cloud segments $\mathbf{S}_{i},i=0,1,2,\cdots$, estimate the corresponding linear segments $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ in the continuous trajectory $\mathbf{T}(t)$. Each linear segment $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ is solved through point-to-plane registration, constant velocity constraint. In summary, we solve the following problem to get the control nodes $\mathbf{X}=(\mathbf{T}_{i},\mathbf{T}_{i+1},\cdots,\mathbf{T}_{i+M})$ in the sliding window, $M$ is the capacity of sliding window:

### v3

@@ Abstract
The motion distortion in LiDAR scans caused by the robot’s aggressive motion and environmental terrain features significantly impacts the positioning and mapping performance of 3D LiDAR odometry. Existing distortion correction solutions struggle to balance computational complexity and accuracy. In this letter, we propose an Adaptive Temporal Interval-based Continuous-Time LiDAR-only Odometry (ATI-CTLO), which is based on straightforward and efficient linear interpolation. Our method can flexibly adjust the temporal intervals between control nodes according to the motion dynamics and environmental degeneracy. This adaptability enhances performance across various motion states and improves the algorithms robustness in degenerate, particularly feature-sparse, environments. We validated our method’s effectiveness on multiple datasets across different platforms, achieving comparable accuracy to state-of-the-art LiDAR-only odometry methods. Notably, in situations involving aggressive motion and sparse features, our method outperforms existing LiDAR-only methods.
@@ I Introduction
LiDAR odometry, recognized for its precise measurement capabilities and adaptability to varying light and weather conditions [1], has been extensively utilized in robotics and autonomous driving, demonstrating remarkable progress over the last decade. Nonetheless, LiDAR odometry faces significant practical challenges, particularly the motion distortion in point clouds, which critically affects localization accuracy and map-building quality. There are several sources of motion distortion, including robot’s movement during data acquisition, uneven terrain that causes the robot to oscillate, and environmental disturbances such as air currents affecting unmanned aerial vehicles (UAVs). Fig. 1 illustrates the motion distortion resulting from aggressive turning during data acquisition and the mapping result of our method.
@@
Most current LiDAR odometry methods are based on discrete-time methods, which assume LiDAR remains stationary during data acquisition and handle LiDAR distortion as part of preprocessing, aligning all points in the point cloud to a common reference frame based on constant velocity assumption [2, 3]. This method, while straightforward and efficient, is inadequate for more aggressive motions (As Fig. 1b shows,  [3] remains distorted). In LiDAR-inertial odometry (LIO) systems, Inertial Measurement Units (IMUs) are used to directly measure the poses of LiDAR points to eliminate distortions [4, 5]. However, this requires precise calibration and is highly susceptible to platform vibrations, such as mechanical vibrations of quadruped robots [6], and is limited by the IMU’s measurement range in extreme motion conditions [7].
@@
Figure 1: Motion distortion correction. (a): A single frame of the point cloud is distorted due to aggressive turning. (b): Correction distortion under constant velocity assumption (The mapping result of KISS-ICP [3]). (c): Our method, utilizing dynamic temporal interval control, ensures more accurate mapping.
@@
In contrast, continuous-time methods consider the LiDAR to be moving during data acquisition and model the trajectory as a function of time, allowing for pose queries at any given time, offering a novel approach to correcting motion distortion. Current models often assume constant velocity [8, 9, 10] or higher-order modeling [11, 12, 13]. However, higher-order models increase computational complexity, while lower-order models suffer from reduced precision. One solution to this problem relies on the assumption of motion continuity, using lower-order modeling but with smaller temporal intervals between the control nodes [8, 14]. Smaller temporal interval allows the constant velocity motion prior to accommodate aggressive movements, but also introduces a new challenge. The small temporal intervals in continuous-time methods are often shorter than the duration of a single LiDAR frame, making it essential to segment the point cloud temporally. This segmentation reduces constraints between control nodes, leading to potential algorithm degeneracy or failure in feature-sparse environments.
@@
To address the aforementioned challenges, we propose a continuous-time LiDAR-only method, the main contributions of our method are as follows:
@@
  • To address the issue of degeneracy due to cloud segmentation, we introduce a degeneracy management method for continuous-time LiDAR odometry. By categorizing the level of degeneracy based on the strength of environmental information constraints on the pose, we implement targeted treatments for each level. This approach significantly enhances the robustness of odometry in degenerate environments.
  • Experiments conducted on various platforms, including unmanned ground vehicles (UGVs), UAVs, handheld devices, and quadruped robots, have demonstrated the effectiveness of our method. Under highly dynamic conditions, our algorithm achieves impressive improvement in accuracy compared to state-of-the-art (SOTA) methods. Additionally, in environments with sparse features, our method exhibits remarkable stability and outperforms all existing LiDAR-Only odometry methods.
@@ II-A Discrete-time and Continuous-time LiDAR Odometry
3D LiDAR odometry technology primarily estimates pose through point cloud registration. The classic method, Iterative Closest Point (ICP) algorithm [15] is widely used. For instance, KISS-ICP [3] proposed a lightweight ICP method that dynamically adjusts the data association threshold, adapting to multiple scenarios with just one parameter file. Feature-based methods achieve rapid and precise registration by extracting significant points from the point cloud. LOAM [16] is a representative of this approach, inspiring subsequent advancements like LeGO-LOAM [17], which incorporated ground optimization, and F-LOAM [2], which increased efficiency without sacrificing accuracy by adopting a scan-to-map method.
However, these methods rely on discrete-time representation, which inherently leads to LiDAR distortion issues. The common solutions involve using a constant velocity assumption or leveraging IMU pose estimation [4, 5]. Continuous-time trajectory methods offer an alternative approach to removing motion distortions in point clouds, which allows query poses at any given time. Over the past decade, these techniques have made significant advancements in real-time performance and accuracy. Continuous-time methods can be divided into two main categories: nonparametric and parametric.
The nonparametric methods model the continuous-time trajectory as Gaussian Process (GP) [18, 19, 20], the GP prior smooths terms between discrete-time nodes. For parametric methods, B-spline is widely utilized to model the trajectory due to the locality and smoothness characteristics[21]. Thanks to the efficient derivative computation of cumulative B-splines[22], numerous high-performing B-spline-based continuous-time LiDAR odometry[11, 23] approaches have been developed. These approaches excel in handling point cloud motion distortion and improving the smoothness of the trajectory.
@@ II-B Degeneracy Detection
LiDAR odometry is highly susceptible to degeneracy in environments with sparse or repetitive geometric features. Continuous-time LiDAR odometry, which uses smaller time intervals for more accurate trajectory fitting, exacerbates the risk of degeneracy due to the reduced number of LiDAR points between adjacent control nodes. Hence, degeneracy detection is essential for improving the robustness of LiDAR odometry.
Zhang et al. [25] determined degeneracy by comparing the smallest eigenvalue $\lambda_{min}$ of the Hessian matrix to a predefined threshold, separating the degenerate directions in the state space and optimizing only in well-constrained directions. Hinduja et al. [26] applied this method to ICP, and using the ratio between the largest and smallest eigenvalues as the eigenvalue threshold. Ren et al. [27] suggested that pose updates in the degenerate direction cause oscillations and calculate the covariance matrix of pose updates during optimization, summing the diagonal elements as the degeneracy factor. Nubert et al. [28] introduced a learning-based degeneracy detection method, directly estimating localizability on raw LiDAR scans.
Another approach directly identifies the degenerate directions through the geometric characteristics of the environment. Yang et al. [29] utilized Principal Component Analysis (PCA) to analyze the normal vectors of all points in the point cloud, also using the smallest eigenvalue $\lambda_{min}$ as a criterion for describing degeneracy. Zhen et al. [30] evaluated the strength of the constraint by measuring the sensitivity of measurements w.r.t the robot pose, but the method requires a prior-build map of the environment. X-ICP [31] employed a similar method to [30], and evaluated the contribution of information pairs to the constraints. The method categorized localizability into three levels {localizable, partially localizable, nonlocalizable} and enhanced constraints for partial localizability through resampling the available information.
@@ III-A System Overview
Figure 2: System Overview. Firstly, LiDAR scans segmentation is based on PCA pre-assessment, followed by solving through a sliding window approach (window’s size $M$: 3). Simultaneously, degeneracy management merges cloud segments when degeneracy occurs.
The Fig.2 shows our pipeline. We denote the trajectory as $\mathbf{T}(t)$ in world frame, which is composed of continuous linear segments $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ over temporal interval $[t_{i},t_{i+1}]$, the Lie Group $\mathbf{T}_{i}\in\mathrm{SE(3)}$ represents the pose at timestamp $t_{i}$. For the latest LiDAR input point cloud $\mathbf{L}_{curr}$, the Principal Component Analysis (PCA) technique is first used to evaluate the angle $\boldsymbol{\varphi}$ between the principal directions of $\mathbf{L}_{curr}$ and the previous point cloud $\mathbf{L}_{pre}$. Then the temporal interval $\Delta t$ between control nodes is reduced by a certain factor $\alpha$ ($\Delta t\leftarrow\alpha\Delta t$) once the angle $\boldsymbol{\varphi}$ exceeds the predefined threshold $\mathcal{K}_{vec}$ (Section III-B). After $\Delta t$ has been confirmed, the points in $\mathbf{L}_{curr}$ are divided into cloud segments $\{\mathbf{S}_{i+1},\mathbf{S}_{i+2},\cdots,\mathbf{S}_{i+5}\}$ according to $\Delta t$, each cloud segment $\mathbf{S}_{i}$ corresponds to a continuous linear segment $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$. A sliding window approach is utilized to optimize multiple segments $\{\mathbf{S}_{i},\mathbf{S}_{i+1},\cdots,\mathbf{S}_{i+M-1}\}$ at once ($M$ is the capacity of the sliding window). Once the optimization is finished, the sliding window moves backward to include new linear segments, while the oldest segment is added to the local map (Section III-C). Notably, during the iterative optimization, the degeneracy detection module is applied to determine if the temporal interval should be increased, which is then achieved by merging linear segments (Section III-D).
@@ III-B PCA Pre-assessment to Reduce Temporal Interval
Figure 3: PCA Pre-assessment. A sudden change in the eigenvectors $\boldsymbol{v}_{i}$ indicates an aggressive turn in motion. In the upper image, the nodes with numbers show significant changes in the PCA direction. However, only the yellow nodes (2, 3, 4) are due to motion turns, while the red nodes (1, 5) result from environmental changes, such as the robot moving between a room and a hallway. The line chart below records the maximum PCA direction change $\boldsymbol{\varphi}_{max}$ (blue) and the maximum eigenvalue change $\mathcal{V}_{max}$ (yellow) throughout the entire trajectory. At nodes 1 and 5, there is a significant change in $\mathcal{V}_{max}$, and we filter out erroneous estimates (red nodes) by setting a threshold (red line).
@@
In this letter, the PCA technique is utilized to quickly assess changes in motion states. As shown in Fig. 3, during the robot’s turning process (yellow nodes 2, 3, 4), the principal direction of the point cloud undergoes significant changes (blue). Note that we only concentrate on the changes in the principal direction caused by violent rotations, since constant velocity prior introduces significant errors when such rotations occur.
Considering all $N$ points ${}^{L}\mathbf{p}_{j}$ in $\mathbf{L}_{curr}$, we first compute the covariance matrix $\mathbf{C}\in\mathbb{R}^{3\times 3}$ of all points:
$\mathbf{C}=\sum_{j=1}^{N}(^{L}\mathbf{p}_{j}-\overline{\mathbf{p}})^{\top}(^{L}\mathbf{p}_{j}-\overline{\mathbf{p}}),\quad\overline{\mathbf{p}}=\frac{1}{N}\sum_{j=1}^{N}{{}^{L}\mathbf{p}_{j}}$ (1)
Then Singular Value Decomposition (SVD) is applied to $\mathbf{C}$ to derive the eigenvalues $\lambda_{0}<\lambda_{1}<\lambda_{2}$ and corresponding eigenvectors $\boldsymbol{v}_{0},\boldsymbol{v}_{1},\boldsymbol{v}_{2}$. The eigenvectors represent the principal directions of LiDAR scans, so the angles $\boldsymbol{\varphi}=[\boldsymbol{\varphi}_{1},\boldsymbol{\varphi}_{2},\boldsymbol{\varphi}_{3}]$ between the principal directions $\boldsymbol{v}^{curr},\boldsymbol{v}^{last}$ of adjacent LiDAR scans indicate the dynamics of motion.
@@
However, the principal directions derived from PCA can abruptly change due to changes in the environment. During our experiments, we observed that these environmental changes also lead to variations in the eigenvalues, as illustrated in Fig. 3 (red nodes 1, 5). Leveraging this observation, we filter out changes in the PCA principal directions resulting from environmental changes:
@@
Only when the max change in eigenvalues $\mathcal{V}_{max}<\mathcal{K}_{val}$ do we consider the principal direction change to be caused by motion. In this case, we compare the maximum change in the point cloud’s principal direction $\boldsymbol{\varphi}_{max}=max(\boldsymbol{\varphi}_{final})=max(\boldsymbol{\varphi}_{1},\boldsymbol{\varphi}_{2},\boldsymbol{\varphi}_{3})$. If $\boldsymbol{\varphi}_{max}>\mathcal{K}_{vec}$, it indicates a significant change in the platform’s movement, particularly in orientation, prompting a reduction in the point cloud segment size: $\Delta t\leftarrow\alpha\Delta t$. Alternatively, if $\mathcal{V}_{max}\geq\mathcal{K}_{val}$, this suggests that the principal direction change is due to environmental changes. Under these circumstances, $\boldsymbol{\varphi}_{final}$ is assigned a value of zero, and the point cloud segmentation is performed based on the initial temporal interval $\Delta t$.
III-C Non-Uniform Linear Interpolation Continuous-time LiDAR-Only Odometry Based on Sliding Window
The continuous-time LiDAR odometry addresses the following problem: given the cloud segments $\mathbf{S}_{i},i=0,1,2,\cdots$, estimate the corresponding linear segments $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ in the continuous trajectory $\mathbf{T}(t)$. Each linear segment $[\mathbf{T}_{i},\mathbf{T}_{i+1}]$ is solved through point-to-plane registration, constant velocity constraint. In summary, we solve the following problem to get the control nodes $\mathbf{X}=(\mathbf{T}_{i},\mathbf{T}_{i+1},\cdots,\mathbf{T}_{i+M})$ in the sliding window, $M$ is the capacity of the sliding window: